A photographer wants to mix 3 liters of a 4% acetic acid solution with a 15% solution to get a 9% solution. how many liters of 15% solution should be used to obtain the 9% solution?
Answer by stanbon(75887)

A photographer wants to mix 3 liters of a 4% acetic acid solution with a 15% solution to get a 9% solution. how many liters of 15% solution should be used to obtain the 9% solution?
---------
Equation:
acid + acid = acid
0.04*3 + 0.15x = 0.09(3+x)
=================
4*3 + 15x = 9*3 + 9x
6x = 5*3
x = (1/2)5 = 2.5 liters (amt. of 15% solution needed)
===========
Cheers,
Stan H.
=================